26.The K, for the following reaction A B is 0.020 at 25 °C. a. Write the equilibrium constant expression K, for this reaction a and b please


c. Based on the ICE table completed in partb and the reaction quotient at the beginning one or two sentences (not more than t c,d, and 3 too please
Answer #1

20. 0 K = 0.02= 12/03 Unit : - tom | E 0.05-2 c m ____ en The reaction will proceed towarde | the product side to reach equilconcentration of products concentration of reactants - = 0.06.* 0.02 = 0.05- - 2000-es Do, 001-0.02%= 2 = 2 = 0.00 = 0,00098=